# GANQ: GPU-Adaptive Non-Uniform Quantization for Large Language Models

### v2

@@ 4.2 Main Results
Table 2: WikiText2 perplexity ($\downarrow$) of quantized models under 4-bit and 3-bit. GANQ outperforms state-of-the-art methods.
@@
Weight-only Quantization. The results in Table 2 present the WikiText2 perplexity of various quantized models under 4-bit and 3-bit configurations across different model sizes (with additional perplexity results on the C4 and PTB datasets in Appendix B). As shown, GANQ consistently outperforms baseline methods such as RTN, GPTQ, and OminiQuant across all configurations. For 4-bit quantization, GANQ achieves the lowest perplexity across both OPT and LLaMA models, with notable improvements. Remarkably, on OPT-2.7B, GANQ’s perplexity (12.33) even outperforms the full-precision FP16 model (12.47). GANQ also demonstrates strong performance with 3-bit quantization, maintaining competitive perplexity reductions across model sizes. For example, on OPT-6.7B, GANQ’s perplexity is 11.39, compared to 15.11 for GPTQ and 13.47 for OminiQuant. These results underscore GANQ’s effectiveness in both 4-bit and 3-bit quantization, achieving substantial perplexity reductions across various model scales. The “–” in Table 2 indicates that OmniQuant cannot quantize LLaMA-3-8B on a single RTX 4090 GPU due to memory constraints or the unavailability of the pre-quantized model.
@@
Weight-only Quantization with Outlier Handling. To mitigate outlier impact, methods like RTN, GPTQ, AWQ, and OminiQuant divide per-channel distributions into smaller blocks (typically of size 128). SqueezeLLM retains a small percentage of outliers (e.g., 0.5%) and a fixed number of full rows (default: 10). GANQ can integrate seamlessly with SqueezeLLM’s outlier handling mechanism. We evaluate this integration through experiments. Due to memory constraints, OmniQuant cannot quantize LLaMA-3-8B, and SqueezeLLM is limited to models up to 2.7B. We use results directly from their paper for these cases if available. Additionally, SqueezeLLM’s current code does not support OPT-350M. For GANQ, we retain 0.5% outliers for all OPT models and LLaMA-3-8B. Additionally, we retain 10 full rows for LLaMA-7B and LLaMA-2-7B to ensure a fair comparison with SqueezeLLM
As shown in Table 4, GANQ (indicating GANQ integrated with outlier handling) outperforms other baselines. Furthermore, when retaining only 0.5% of outliers for LLaMA-7B and LLaMA-2-7B, we observe the following results: LLaMA-7B (5.78 for 4-bit, 6.20 for 3-bit), LLaMA-2-7B (5.60 for 4-bit, 6.10 for 3-bit), which still outperform all other methods, except for SqueezeLLM.
Table 4: WikiText2 perplexity ($\downarrow$) of quantized models under 4-bit and 3-bit. GANQ outperforms state-of-the-art methods.
@@
Table 5: Comparison CUDA time (s), speedup ($\uparrow$), and peak memory (GB) ($\downarrow$) for uniform and LUT-based quantization methods on OPT-6.7B and LLaMA-7B models.
@@ 4.3 Profiling
We present the CUDA time, speedup, and peak GPU memory usage of GANQ in Table 5, measured on a single NVIDIA RTX 4090 GPU across different configurations when generating 1024 tokens. GANQ achieves up to a 2.57$\times$ speedup compared to the FP16 baseline, with peak memory usage reduced to 4.10 GB for OPT-6.7B and 3.30 GB for LLaMA-7B at 3-bit quantization. Furthermore, the lack of native support for mpGEMM in uniform quantization methods (e.g., GPTQ) significantly slows down the overall inference process. In contrast, GANQ incurs only a slight increase in memory usage compared to uniform quantization, underscoring that the overhead of LUT-based dequantization is minimal, especially when weighed against the significant improvements in perplexity and latency. Note that the GPTQ-for-LLaMA (GPTQ-for-LLaMa, 2023) package does not currently support 3-bit inference, so no 3-bit results are reported for GPTQ. We expect similar performance to the 4-bit quantization results.
@@ 4.4 Quantization Cost
Among the evaluated methods, RTN, GPTQ, and AWQ are the most efficient in GPU memory usage and quantization time, due to their layer-wise heuristic approach. However, they trade off model performance. In contrast, OmniQuant and SqueezeLLM require gradient information, leading to higher memory demands. OmniQuant can quantize 7B models on a single RTX 4090 GPU but fails for 8B models and takes over 3 hours for 7B quantization. SqueezeLLM, which requires global gradients, can only quantize models up to 2.7B. GANQ leverages GPU-adaptive, parallel row-wise computation to quantize 7B models in approximately 1 hour for $K=10$. Considering both model performance and quantization cost, GANQ is an effective solution.
@@ 5 Conclusion
Appendix A Outlier Extraction Method
@@
Appendix B Additional Results
@@
The results in Table 6 present the C4 perplexity of various quantized models under 4-bit and 3-bit configurations across different model sizes. As shown, GANQ outperforms baseline methods such as RTN, GPTQ, and OminiQuant across all configurations. For 4-bit quantization, GANQ achieves the lowest perplexity across both OPT and LLaMA models, with notable improvements. GANQ also demonstrates strong performance with 3-bit quantization, maintaining competitive perplexity reductions across model sizes. On OPT-6.7B, GANQ’s perplexity is 13.68, compared to 17.00 for GPTQ and 15.51 for OminiQuant. These results highlight GANQ’s effectiveness in both 4-bit and 3-bit quantization, achieving substantial perplexity reductions across a wide range of model scales. The symbol ’–’ in Table 6 indicates that OmniQuant cannot quantize LLaMA-3-8B on a single RTX 4090 GPU due to memory constraints, or the quantized model is unavailable in their model zoo.
@@
The results in Table 7 present the PTB perplexity of various quantized models under 4-bit and 3-bit configurations across different model sizes. As shown, GANQ outperforms baseline methods such as RTN, GPTQ, and OminiQuant across all configurations. For 4-bit quantization, GANQ achieves the lowest perplexity across OPT models with notable improvements. GANQ also demonstrates strong performance with 3-bit quantization, maintaining competitive perplexity reductions across model sizes. On OPT-6.7B, GANQ’s perplexity is 16.91, compared to 21.63 for GPTQ and 21.52 for OminiQuant. These results highlight GANQ’s effectiveness in both 4-bit and 3-bit quantization, achieving substantial perplexity reductions across a range of model sizes.

### v3

@@ 4.2 Main Results
Table 2: WikiText-2 perplexity ($\downarrow$) of quantized models under 4-bit and 3-bit. GANQ outperforms state-of-the-art methods.
@@
Table 4: Performance comparison of 4-bit quantized LLaMA-3.2 1B-Instruct and 3B-Instruct models on LongBench ($\uparrow$) and GSM8K ($\uparrow$) datasets.
Method 1B-Instruct 3B-Instruct
LongBench GSM8K (%) LongBench GSM8K (%)
FP16 11.5 32.90 12.7 64.97
RTN 0.2 4.17 12.1 37.68
GPTQ 11.14
OmniQuant 8.9 16.53 11.5 54.74
GANQ 11.7 27.75 12.5 60.50
Weight-only Quantization. The results in Table 2 present the WikiText-2 perplexity of quantized OPT, LLaMA, LLaMA-2, and LLaMA-3 models under 4-bit and 3-bit configurations across different model sizes (with additional perplexity results on the C4 and PTB datasets as well as results of quantized LLaMA-3.2 models in Appendix C). As shown, GANQ consistently outperforms baseline methods such as RTN, GPTQ, and OmniQuant across all configurations. For 4-bit quantization, GANQ achieves the lowest perplexity across both OPT and LLaMA models, with notable improvements. Remarkably, on OPT-2.7B, GANQ’s perplexity (12.33) even outperforms the full-precision FP16 model (12.47). GANQ also demonstrates strong performance with 3-bit quantization, maintaining competitive perplexity reductions across model sizes. For example, on OPT-6.7B, GANQ’s perplexity is 11.39, compared to 15.11 for GPTQ and 13.47 for OmniQuant. These results underscore GANQ’s effectiveness in both 4-bit and 3-bit quantization, achieving substantial perplexity reductions across various model scales. The “–” in Table 2 indicates that OmniQuant cannot quantize LLaMA-3-8B on a single NVIDIA RTX 4090 GPU due to memory constraints or the unavailability of the pre-quantized model.
@@
Table 4 presents a performance comparison of 4-bit quantized LLaMA-3.2 1B-Instruct and 3B-Instruct models on two tasks: LongBench, which evaluates long-context capabilities, and GSM8K, which assesses reasoning-intensive tasks. Both tasks are evaluated in zero-shot settings without chain-of-thought prompting. GANQ consistently outperforms baseline methods on both tasks, highlighting its robustness. In contrast, RTN exhibits extreme sensitivity to model scale, collapsing on LLaMA-3.2-1B-Instruct. Similarly, GPTQ achieves near-zero GSM8K accuracy on LLaMA-3.2-3B-Instruct, consistent with the perplexity results reported in Table 10, and encounters LongBench execution errors (i.e., skip_special_tokens failures) despite identical experimental setups. Compared to RTN and GPTQ, OmniQuant leverages learnable quantization parameters, leading to relatively stable performance. However, it still falls short of GANQ’s results. These findings underscore GANQ’s ability to deliver superior performance across diverse tasks.
Weight-only Quantization with Outlier Handling. To mitigate outlier impact, methods like RTN, GPTQ, AWQ, and OmniQuant divide per-channel distributions into smaller blocks (typically of size 128). SqueezeLLM retains a small percentage of outliers (e.g., 0.5%) and a fixed number of full rows (default: 10). GANQ can integrate seamlessly with SqueezeLLM’s outlier handling mechanism. We evaluate this integration through experiments. Due to memory constraints, OmniQuant cannot quantize LLaMA-3-8B, and SqueezeLLM is limited to models up to 2.7B. We use results directly from their paper for these cases if available. Additionally, SqueezeLLM’s current code does not support OPT-350M. For GANQ, we retain 0.5% outliers for all OPT models and LLaMA-3-8B. Additionally, we retain 10 full rows for LLaMA-7B and LLaMA-2-7B to ensure a fair comparison with SqueezeLLM.
As shown in Table 5, GANQ (indicating GANQ integrated with outlier handling) outperforms other baselines. Furthermore, when retaining only 0.5% of outliers for LLaMA-7B and LLaMA-2-7B, we observe the following results: LLaMA-7B (5.78 for 4-bit, 6.20 for 3-bit), LLaMA-2-7B (5.60 for 4-bit, 6.10 for 3-bit), which still outperform all other methods except for SqueezeLLM.
Table 5: WikiText-2 perplexity ($\downarrow$) of quantized models under 4-bit and 3-bit. GANQ outperforms state-of-the-art methods.
@@
Table 6: CUDA time (s), speedup ($\uparrow$), and peak memory (GB) ($\downarrow$) for GANQ and GANQ using the LUT-based inference kernel from (Kim et al., 2024) on OPT-6.7B and LLaMA-7B models.
@@ 4.3 Profiling
Table 6 reports CUDA time, speedup, and peak memory usage for GANQ and GANQ, measured on a single NVIDIA RTX 4090 GPU while generating 1,024 tokens with a batch size of 1 on OPT-6.7B and LLaMA-7B. GANQ achieves up to a 2.57$\times$ speedup over the FP16 baseline, with peak memory usage reduced to 4.10 GB for OPT-6.7B and 3.30 GB for LLaMA-7B under 3-bit quantization. While GANQ provides improved quantization accuracy through outlier handling, it incurs slightly higher inference latency due to additional sparse matrix operations. Therefore, the choice between GANQ and GANQ should be guided by the desired trade-off between accuracy and runtime efficiency.
The above results demonstrate that GANQ is fully compatible with existing LUT-based inference kernels (Kim et al., 2024). Moreover, GANQ stands to benefit from ongoing engineering advancements in this class of kernels, such as those proposed in (Guo et al., 2024; Mo et al., 2024), which can further enhance implementation efficiency and overall inference performance.
@@ 4.4 Quantization Cost
Among the evaluated methods, RTN, GPTQ, and AWQ are the most efficient in GPU memory usage and quantization time, due to their layer-wise heuristic approach. However, they trade off model performance. In contrast, OmniQuant and SqueezeLLM require gradient information, leading to higher memory demands. OmniQuant can quantize 7B models on a single NVIDIA RTX 4090 GPU but fails for 8B models and takes over 3 hours for 7B quantization. SqueezeLLM, which requires global gradients, can only quantize models up to 2.7B. GANQ leverages GPU-adaptive, parallel row-wise computation to quantize 7B models in approximately 1 hour for $K=10$. Considering both model performance and quantization cost, GANQ is an effective solution.
@@ 5 Conclusion
Acknowledgments
The work of X. Yuan was supported by the GRF RGC (No. 17309824).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Ensuring Positive Definiteness via Diagonal Dominance
In some cases, such as the fc2 layer of OPT models, the matrix $\mathbf{X}\mathbf{X}^{\top}$ may fail to be positive definite. While this is uncommon in our experiments, it is important to handle such situations before performing Cholesky decomposition in Algorithm 1, which requires a positive definite matrix. To ensure this, we enforce diagonal dominance via a small adjustment to the matrix.
Specifically, a symmetric matrix $\mathbf{A}$ is guaranteed to be positive definite if it is diagonally dominant with strictly positive diagonal entries, meaning $|a_{ii}|\geq\sum_{j\neq i}|a_{ij}|$ for all $i$. To utilize this property, let us denote $\bm{\Sigma}=\mathbf{X}\mathbf{X}^{\top}$, where diagonal elements satisfy $\bm{\Sigma}_{ii}\geq 0$. We compute an adaptive offset vector $\bm{\delta}$, where each component $\bm{\delta}_{i}$ is calculated as:
[…]
where $\bm{\Sigma}_{i,j}$ denotes the element at the $i$-th row and $j$-th column of $\bm{\Sigma}$. Subsequently, we obtain a diagonally dominant matrix by adding this adaptive diagonal offset and perform the Cholesky decomposition:
$\mathbf{L}=\mathrm{Cholesky}\left(\bm{\Sigma}+\mathrm{Diag}(\bm{\delta})\right),$ (24)
where $\mathrm{Diag}(\bm{\delta})$ generates a diagonal matrix from the vector $\bm{\delta}$, and $\mathbf{L}$ is the resulting lower-triangular Cholesky factor. This adaptive preconditioning procedure efficiently ensures numerical stability and positive definiteness, facilitating robust computation without manual parameter adjustments.
To evaluate sensitivity to preconditioning approaches (the standard fixed $\lambda$ illustrated in Remark 3.1 versus our adaptive method), we conduct 4-bit quantization experiments on OPT-125M and report the resulting perplexity on WikiText-2 in Table 7.
Table 7: WikiText-2 perplexity ($\downarrow$) of 4-bit quantized OPT-125M under different preconditioning strategies.
     $\mathbf{X}\mathbf{X}^{\top}+\lambda\mathbf{I}$      Method in (23) – (24)
     $\lambda=0.5$      $\lambda=1.0$      $\lambda=10.0$      $\lambda=40.0$      $\lambda=100.0$
     29.14      29.04      28.98      29.05      29.09      28.58
The results show that quantization accuracy of GANQ is largely insensitive to the choice of preconditioning. The adaptive method achieve the best result, with fixed $\lambda$ yielding similar results. All results outperforms baselines, confirming the robustness.
Appendix B Outlier Extraction Method
@@
Appendix C Additional Results
The results in Table 8 present the C4 perplexity of various quantized models under 4-bit and 3-bit configurations across different model sizes. As shown, GANQ outperforms baseline methods such as RTN, GPTQ, and OmniQuant across all configurations. For 4-bit quantization, GANQ achieves the lowest perplexity across both OPT and LLaMA models, with notable improvements. GANQ also demonstrates strong performance with 3-bit quantization, maintaining competitive perplexity reductions across model sizes. On OPT-6.7B, GANQ’s perplexity is 13.68, compared to 17.00 for GPTQ and 15.51 for OmniQuant. These results highlight GANQ’s effectiveness in both 4-bit and 3-bit quantization, achieving substantial perplexity reductions across a wide range of model scales. The symbol “–” in Table 8 indicates that OmniQuant cannot quantize LLaMA-3-8B on a single NVIDIA RTX 4090 GPU due to memory constraints, or the quantized model is unavailable in their model zoo.
@@ Appendix C Additional Results
The results in Table 9 present the PTB perplexity of various quantized models under 4-bit and 3-bit configurations across different model sizes. As shown, GANQ outperforms baseline methods such as RTN, GPTQ, and OmniQuant across all configurations. For 4-bit quantization, GANQ achieves the lowest perplexity across OPT models with notable improvements. GANQ also demonstrates strong performance with 3-bit quantization, maintaining competitive perplexity reductions across model sizes. On OPT-6.7B, GANQ’s perplexity is 16.91, compared to 21.63 for GPTQ and 21.52 for OmniQuant. These results highlight GANQ’s effectiveness in both 4-bit and 3-bit quantization, achieving substantial perplexity reductions across a range of model sizes.
@@
Table 10: Perplexity ($\downarrow$) of quantized LLaMA-3.2 models on WikiText2 and C4. GANQ outperforms state-of-the-art methods.
Dataset Method Bit-width 1B 3B 1B-Instruct 3B-Instruct
WikiText-2 FP16 16 9.76 7.81 13.16 11.05
RTN 4 18.08 10.53 22.91 15.58
GPTQ 4 24.07 6.0e3 18.40 4.5e3
OmniQuant 4 12.90 8.87 16.31 12.33
GANQ 4 10.78 8.35 14.36 11.99
RTN 3 2.6e3 4.8e2 7.0e3 5.7e2
GPTQ 3 1.68e2 6.3e3 1.1e2 3.3e3
OmniQuant 3 4.3e2 14.82 34.89 19.55
GANQ 3 15.91 10.85 21.32 15.84
C4 FP16 16 14.02 11.34 21.30 16.50
RTN 4 27.98 15.89 35.23 22.15
GPTQ 4 41.27 2.1e3 27.99 3.2e3
OmniQuant 4 18.34 13.08 25.20 18.02
GANQ 4 15.53 12.21 22.81 17.35
RTN 3 2.9e3 4.1e2 6.7e3 5.3e2
GPTQ 3 1.5e2 1.6e3 1.2e2 2.0e3
OmniQuant 3 3.9e2 19.96 41.43 24.80
GANQ 3 22.09 15.36 29.18 21.30
Table 10 presents the perplexity of quantized LLaMA-3.2 models on WikiText-2 and C4 under 4-bit and 3-bit configurations. GANQ consistently outperforms baseline methods (RTN, GPTQ, OmniQuant) across all model variants and bit-widths. Our systematic optimization framework for layer-wise non-uniform quantization allows GANQ to adapt effectively to varying weight distributions, unlike baseline methods, which demonstrate greater sensitivity and instability, especially at 3-bit quantization.